# Spectral Editing of Activations for Large Language Model Alignment

### v2

@@ 2.4 Editing Activations during Inference
The vectors $\overline{\mathbf{z}}^{(t+)}_{\ell},\overline{\mathbf{z}}^{(t-)}_{\ell}$ are the activations after editing negatively and positively in the $\ell$-th layer. These two vectors are merged together to get the final edited activation vectors as follows, where $i$ ranges over the coordinates of the vectors: $\overline{{z}}^{(t)}_{\ell,i}=(\overline{{z}}^{(t+)}_{\ell,i}+\overline{{z}}^{(t-)}_{\ell,i})\times\displaystyle\frac{\sqrt{\sum_{t=1}^{T}(z_{\ell,i}^{(t)})^{2}}}{\sqrt{\sum_{t=1}^{T}(z^{(t+)}_{\ell,i}+z^{(t-)}_{\ell,i})^{2}}}$.
@@ 3.1 Truthfulness
Datasets. We evaluate all compared methods on TruthfulQA [22], which consists of 817 questions in 38 subcategories, each paired with a single best answer, and multiple correct/incorrect answers. TruthfulQA contains two evaluation protocols: multiple-choice question answering, and generation. We mainly use the first to ensure the comparability with previous work [20, 21, 6]. We show an example of a TruthfulQA data instance together with the demonstrations we used in Appendix H.
@@
Evaluation Metrics. Following [40], we adopt the multiple-choice metrics (MC-1/2) from TruthfulQA. MC1 assesses whether the model allocates the highest predicted likelihood to the best answer, while MC2 evaluates whether the normalised likelihood of all correct answers surpasses the incorrect ones.
Baselines. We compare SEA against several baselines: 1) In-Context Learning (ICL): ICL shows that LLMs can learn from demonstrations in the prompt. Here, we test if LLMs can learn to generate truthful responses from the positive demonstrations. 2) LoRA Fine-tuning (LoRA-FT; Hu et al. 14): we use the same training data for SEA to construct an Alpaca-style [35] instruction-tuning dataset, and then we further fine-tune the LLM on this dataset with LoRA [14]. 3) Inference-time Intervention (ITI; Li et al. 20): ITI trains a shifting module to capture truthfulness and then applies it to LLM’s activations during inference. 4) DoLa [6] attempts to improve the model’s factuality by contrasting the model’s predicted logits based on various layers’ activations. 5) Contrastive Decoding (CD; Li et al. 21) manipulates the model’s predicted logits by penalising the ones similar to a smaller model. 6) Induce-then-Contrast Decoding (ICD; Zhang et al. 40) follows the intuition of CD, but ICD replaces the small model with an induced hallucinated model. We use a prompt-based induced hallucinated model here for a fair comparison.
@@ 3.2 Bias
Dataset. We assess the model bias on the Bias Benchmark for QA (BBQ; Parrish et al. 25), which is widely adopted in LLM evaluation [17, 29, 2]. BBQ formulates bias evaluation as a question-answering task, allowing us to easily construct the paired positive and negative demonstrations. BBQ contains 29,246 questions covering 11 diverse types of common bias. We randomly sampled 5,246 questions for evaluation, and the rest were used for training and validation purposes. We use the disambiguated version for a more comprehensive evaluation, providing the model with sufficiently informative contexts. This approach enables us to assess whether the model biases influence its selection of a correct answer choice. See Appendix H for example demonstrations.
@@ 4.1 Truthfulness Evaluation
Main results. Table 1 illustrates the results of various inference-only techniques aimed at boosting the performance of a 7B LLaMA-2 model on TruthfulQA. We observe a significant enhancement of the base LLaMA-2 model when further trained with the conversation-style alignment (LLaMA-2-Chat), as described in [37]. On the other hand, we do not observe an improvement with LoRA, which indicates the difficulty in improving a well-trained model with LoRA-style fine-tuning.
@@
Method MC1 MC2 TInf. TTrain
ICL (LLaMA-2) 28.39 43.42 - -
ICL (LLaMA-2-Chat) 36.96 54.68 4.90 -
LoRA-FT (LLaMA-2-Chat; $N=1000$) 35.74 54.61 5.16 299
LoRA-FT (LLaMA-2-Chat; $N=2000$) 35.01 54.24 5.04 1190
ITI 37.01 54.66 5.82 -
DoLA 32.97 60.84 5.60 -
CD (13B-Chat vs. 7B-Chat) 28.15 54.87 - -
ICD (Prompt-version) 37.87 57.77 9.67 -
SEA ($N=1000,K=99\%,L=4$) 38.31 55.27 5.08 140
SEA ($N=2000,K=99.8\%,L=21$) 39.41 57.15 5.93 152
The best result is from our proposed SEA method, which outperforms ICL by 2.45 and 2.47 points for MC1 and MC2, respectively. When compared to alternative approaches, SEA achieves the highest MC1 score while incurring the minimal sacrifice in inference speed when we only modify the last four layers. SEA has only a 3.67% increase in inference speed, contrasting with the larger increases of 18.78%, 14.29%, and 97.34% for ITI, DoLA, and ICD, respectively. We also highlight the training efficiency, i.e., computing editing projections, of SEA compared to gradient-based optimisation methods (e.g., LoRA), an advantage that becomes more significant with an increasing number of demonstrations.
@@
Our second group of analysis is to confirm the advantage of using feature normalisation. We alter the combination of positively and negatively edited activations by simply averaging each neuron’s activation between the twos. We note a decrease in MC1 from 39.41 to 35.86, affirming the impact of our proposed normalisation technique. One potential explanation for this decline in performance is that basic averaging might disrupt the correlation between activations, thereby hindering subsequent layers of the model from processing the edited activations normally.
@@ 4.2 Bias Evaluation
Methods AUBSSR%
ICL 43.02 19.6 5.8 52.0
LoRA-FT 44.72 27.0 4.3 40.4
SEA
[…]
Sq. Exp. 56.17 8.10 2.0 42.6
Tanh 48.06 11.3 7.3 51.8
ELU 46.57 14.3 8.1 52.6
[…]
Table 3: BBQ performance (in terms of accuracy, A%, unknown-answer rate, Uand stereotypical response rate, SR%) after applying ICL, Linear SEA and non-linear SEA ($\Phi$-SEA) on six open-source LLMs. We highlight the improved and worsened metrics, respectively.
LLaMA-2-7B AUSRLLaMA-2-13B AUSRLLaMA-2-70B AUSR%
ICL 43.0 19.6 52.0 ICL 47.1 18.1 47.6 ICL 45.86 17.5 49.1
Linear SEA 43.8 18.5 51.9 Linear SEA 47.3 17.9 47.1 Linear SEA 45.9 17.7 49.08
$\Phi$-SEA 56.2 8.1 42.6 $\Phi$-SEA 54.6 12.3 44.6 $\Phi$-SEA 58.03 13.4 39.6
Gemma-it-2B AUSRGemma-it-7B AUSRMistral-7B AUSR%
ICL 41.8 20.4 52.5 ICL 44.4 30.1 43.6 ICL 94.6 0.8 4.2
Linear SEA 41.9 19.3 52.7 Linear SEA 44.2 31.2 44.0 Linear SEA 94.8 0.7 4.3
$\Phi$-SEA 44.5 15.7 52.6 $\Phi$-SEA 48.1 30.3 40.6 $\Phi$-SEA 95.7 0.5 3.9
@@
SEA Can be Generalised to Various LLMs. We show the results of applying ICL, Linear SEA and $\Phi$-SEA on six LLMs with different model families and sizes in Table 3. Specifically, we test them on the LLaMA-2 family [37], Gemma family [36] and Mistral 7B [16], which represent the state-of-the-art open-source LLMs. We observe a general trend across different LLMs that linear SEA can marginally improve accuracy, but the other two metrics are mixed. $\Phi$-SEA shows promising performance and can gain increased accuracy, lower unknown-answer rate and stereotypical response rate across all LLMs, except a negligible higher stereotypical response rate for Gemma-it-2B.
@@ 4.4 Post-Editing Performance on Control Tasks
Table 4: Performance of LLaMA-2-Chat-7B and its three SEA edited models for truthfulness and fairness on six control tasks covering multi-task ability, commonsense question answering, and mathematical ability. Details of evaluation are provided in Appendix D.4.
@@
w/ Phi-SEA-Fair 51.93 $\downarrow$ 5.85 14.90 $\downarrow$ 7.92 20.92 $\downarrow$ 1.52 30.05 $\downarrow$ 1.64 45.30 $\downarrow$ 1.15 56.38 $\uparrow$ 5.21
We then leverage six additional control tasks to analyse the editing effects of SEA on other LLM capabilities in Figure 4. These tasks include MMLU [13], which serves as the most general evaluation benchmark. We then use HellaSwag [18] and Natural Questions [39] to assess the model’s commonsense reasoning and question-answering ability. We use GSM8K [7] and MathQA [1] to verify the model’s ability to solve mathematical tasks. We rely on ToxiGen [12] for assessing the model’s toxicity.
Editing a model’s activations has only a limited impact on other capabilities. We first note that linear editing almost does not hurt the model’s other capabilities (e.g., in commonsense and maths). The non-linear editing causes a small drop in performance in maths tasks, but we observe a more visible decrease in common-sense QA. We attribute this to the limitation mentioned in §2.5, namely that the non-linear projection using feature functions is not theoretically lossless. Qualitative examples in Appendix B also show the high quality and fluency in outputs of SEA-edited models.
The editing effect can be generalised to other similar tasks. Finally, we observe that the post-editing improvements can be generalised across new tasks, provided they share some similarities. For example, we observe a strong improvement of the 7B LLaMA-2-Chat model in toxicity evaluation, after linearly or non-linearly projecting the activations towards less biased directions. We also see improvements on MMLU and GSM8K after applying the truthful editing.
@@ 4.5 Spectral Analysis for the Source of Model Behaviours
We now present an analysis to substantiate why editing the top-$L$ layers proves to be more precise compared to other editing schemes, such as the bottom-$L$ layers. We follow [10, 41, 42] which proposed interpreting the signature as the sum of singular values resulting from an SVD of the cross-covariance for two variables, indicating the degree of correlation between them. We extend this idea to identify which layers’ activations contain the most information regarding the model’s behaviour (e.g., generating hallucinated or biased responses). We introduce the details of calculating the signature value in Appendix A.
@@ 5 Related Work
Modifying the activation of a trained model, thus altering the model’s behaviour [28, 15, 23, 43] or internal knowledge [8], represents a lightweight method to control the model’s generation. Li et al. [20] propose an approach to initially identify which attention heads of an LLM are accountable for generating hallucinations, followed by editing activations toward truthful directions during inference. Another way is extracting latent vectors directly from the trained model and leveraging these vectors to regulate the model’s inference [38, 34, 30]. Recently, Singh et al. [33] demonstrated the efficacy of fitting an optimal transport from negative to positive activations to facilitate effective non-linear editing. Activation editing finds application in decoding as well, either by contrasting activations from various layers [6] or by using a weaker model to edit activations from a stronger model [21, 40].

### v3

@@ 2.4 Editing Activations during Inference
The vectors $\overline{\mathbf{z}}^{(t+)}_{\ell},\overline{\mathbf{z}}^{(t-)}_{\ell}$ are the activations after editing negatively and positively in the $\ell$-th layer. These two vectors are merged together to get the final edited activation vectors as follows, where $i$ ranges over the coordinates of the vectors:
$\overline{{z}}^{(t)}_{\ell,i}=(\overline{{z}}^{(t+)}_{\ell,i}+\overline{{z}}^{(t-)}_{\ell,i})\times\displaystyle\frac{\sqrt{\sum_{t=1}^{T}(z_{\ell,i}^{(t)})^{2}}}{\sqrt{\sum_{t=1}^{T}(z^{(t+)}_{\ell,i}+z^{(t-)}_{\ell,i})^{2}}}.$ (4)
@@ 3.1 Truthfulness
Datasets. We evaluate all compared methods on TruthfulQA [23], which consists of 817 questions in 38 subcategories, each paired with a single best answer, and multiple correct/incorrect answers. TruthfulQA contains two evaluation protocols: multiple-choice question answering, and generation. We mainly use the first to ensure the comparability with previous work [21, 22, 7]. We show an example of a TruthfulQA data instance together with the demonstrations we used in Appendix M.
@@
Evaluation Metrics. Following [45], we conduct the evaluation on both the multiple-choice question answering and generation track. In the first track, we use MC-1/2 from TruthfulQA: MC1 assesses whether the model allocates the highest predicted likelihood to the best answer, while MC2 evaluates whether the normalised likelihood of all correct answers surpasses the incorrect ones. In the generation track, we follow Li et al. [21] to train two separate evaluators, GPT-Truth and GPT-Info. Each of evaluators predicts a score from 00 to $1$ indicating the truthfulness and informativeness of a given response, respectively. Additionally, we report the Info*Truth metric which assesses a response if it is both informative and truthful. We also report the inference and training time for LoRA and SEA for the efficiency comparison between the gradient-based method and ours.
Baselines. We compare SEA against several baselines: 1) In-Context Learning (ICL): ICL shows that LLMs can learn from demonstrations in the prompt. Here, we test if LLMs can learn to generate truthful responses from the positive demonstrations. 2) LoRA Fine-tuning (LoRA-FT; Hu et al. 15): we use the same training data for SEA to construct an Alpaca-style [38] instruction-tuning dataset, and then we further fine-tune the LLM on this dataset with LoRA [15]. 3) Inference-time Intervention (ITI; Li et al. 21): ITI trains a shifting module to capture truthfulness and then applies it to LLM’s activations during inference. 4) DoLa [7] attempts to improve the model’s factuality by contrasting the model’s predicted logits based on various layers’ activations. 5) Contrastive Decoding (CD; Li et al. 22) manipulates the model’s predicted logits by penalising the ones similar to a smaller model. 6) Induce-then-Contrast Decoding (ICD; Zhang et al. 45) follows the intuition of CD, but ICD replaces the small model with an induced hallucinated model. We use a prompt-based induced hallucinated model here for a fair comparison.
@@ 3.2 Bias
[…]
Method MC1 MC2 Info Truth Info*Truth TInf. TTrain
ICL (LLaMA-2) 28.39 43.42 - - - - -
ICL (LLaMA-2-Chat) 36.96 54.68 69.40 47.36 33.29 4.90 -
[…]
$w/$ Best-of-$2$ 76.50 57.03 44.55 - -
$w/$ Best-of-$3$ 80.54 62.30 50.31 - -
[…]
LoRA-FT (LLaMA-2-Chat; $N=2000$) 35.01 54.24 92.41 47.49 42.35 5.04 1190
ITI 37.01 54.66 - - - 5.82 -
DoLA 32.97 60.84 - - - 5.60 -
CD (13B-Chat vs. 7B-Chat) 28.15 54.87 - - - - -
ICD (Prompt-version) 37.87 57.77 - - - 9.67 -
SEA ($N=1000,K=99\%,L=4$) 38.31 55.27 70.38 48.96 35.25 5.08 140
SEA ($N=2000,K=99.8\%,L=21$) 39.41 57.15 68.05 50.67 33.66 5.93 152
[…]
$w/$ Best-of-$2$ 77.72 57.28 44.56 - -
$w/$ Best-of-$3$ 82.01 63.04 51.30 - -
Dataset. We assess the model bias on the Bias Benchmark for QA (BBQ; Parrish et al. 27), which is widely adopted in LLM evaluation [18, 31, 2]. BBQ formulates bias evaluation as a question-answering task, allowing us to easily construct the paired positive and negative demonstrations. BBQ contains 29,246 questions covering 11 diverse types of common bias. We randomly sampled 5,246 questions for evaluation, and the rest were used for training and validation purposes. We use the disambiguated version for a more comprehensive evaluation, providing the model with sufficiently informative contexts. This approach enables us to assess whether the model biases influence its selection of a correct answer choice. See Appendix M for example demonstrations.
@@ 4.1 Truthfulness Evaluation
Methods AUBSSR%
ICL 43.02 19.6 5.8 52.0
LoRA-FT 44.72 27.0 4.3 40.4
SEA
@@
Sq. Exp. 56.17 8.10 2.0 42.6
Tanh 48.06 11.3 7.3 51.8
ELU 46.57 14.3 8.1 52.6
[…]
Main results. Table 3.2 illustrates the results of various inference-only techniques aimed at boosting the performance of a 7B LLaMA-2 model on TruthfulQA. We observe a significant enhancement of the base LLaMA-2 model when further trained with the conversation-style alignment (LLaMA-2-Chat), as described in [40]. On the other hand, we do not observe an improvement with LoRA, which indicates the difficulty in improving a well-trained model with LoRA-style fine-tuning.
In the multiple-choice track, the best result is from our proposed SEA method, which outperforms ICL by 2.45 and 2.47 points for MC1 and MC2, respectively. In the generation track, SEA has the better truthfulness compared with ICL and LoRA. The positive improvement is also observed in the Best-of-N distribution. When compared to alternative approaches, SEA achieves the highest MC1 score while incurring the minimal sacrifice in inference speed when we only modify the last four layers. SEA has only a 3.67% increase in inference speed, contrasting with the larger increases of 18.78%, 14.29%, and 97.34% for ITI, DoLA, and ICD, respectively. We also highlight the training efficiency, i.e., computing editing projections, of SEA compared to gradient-based optimisation methods (e.g., LoRA), an advantage that becomes more significant with an increasing number of demonstrations.
@@
Our second group of analysis is to confirm the advantage of using feature normalisation. We alter the combination of positively and negatively edited activations by simply averaging each neuron’s activation between the two. We note a decrease in MC1 from 39.41 to 35.86, affirming the impact of our proposed normalisation technique. One potential explanation for this decline in performance is that basic averaging might disrupt the correlation between activations, thereby hindering subsequent layers of the model from processing the edited activations normally.
@@
Table 3: BBQ performance (in terms of accuracy, Acc.%, unknown-answer rate, Unk.and stereotypical response rate, SR%) and TruthfulQA performance (in terms of MC1/2) after applying ICL, Linear SEA and non-linear SEA ($\Phi$-SEA) on six open-source LLMs. We highlight the improved and worsened metrics, respectively.
$\Delta$ Model Size LLaMA-2-7B LLaMA-2-13B LLaMA-2-70B
ICL Linear-SEA $\Phi$-SEA ICL Linear-SEA $\Phi$-SEA ICL Linear-SEA $\Phi$-SEA
BBQ Acc.43.0 43.8 ($\uparrow 0.8$) 56.2 ($\uparrow 13.2$) 47.1 47.3 $(\uparrow 0.2)$ 54.6 $(\uparrow 7.5)$ 45.8 45.9 $(\uparrow 0.1)$ 58.0 $(\uparrow 12.2)$
Unk.19.6 18.5 ($\downarrow 1.1$) 8.10 ($\downarrow 11.5$) 18.1 17.9 $(\downarrow 0.2)$ 12.3 $(\downarrow 5.6)$ 17.5 17.7 $(\uparrow 0.2)$ 13.4 $(\downarrow 4.10)$
SR52.0 51.9 ($\downarrow 0.1$) 42.6 ($\downarrow 9.40$) 47.6 47.1 $(\downarrow 0.5)$ 44.6 $(\downarrow 3.0)$ 49.1 49.1 $(=0.0)$ 39.6 $(\downarrow 9.50)$
TruthfulQA MC1 36.9 39.4 ($\uparrow 2.5$) / 37.7 38.0 $(\uparrow 0.3)$ / 37.7 37.8 $(\uparrow 0.1)$ /
MC2 54.6 57.1 ($\uparrow 2.5$) / 55.7 55.6 $(\downarrow 0.1)$ / 59.0 58.9 $(\downarrow 0.1)$ /
$\Delta$ Model Family Gemma-it-2B Gemma-it-7B Mistral-7B
ICL Linear-SEA $\Phi$-SEA ICL Linear-SEA $\Phi$-SEA ICL Linear-SEA $\Phi$-SEA
BBQ Acc.41.8 41.9 $(\uparrow 0.1)$ 44.5 $(\uparrow 2.7)$ 44.4 44.2 $(\downarrow 0.2)$ 48.1 $(\uparrow 3.7)$ 94.6 94.8 $(\uparrow 0.2)$ 95.7 $(\uparrow 1.1)$
Unk.20.4 19.3 $(\downarrow 1.1)$ 15.7 $(\downarrow 4.7)$ 30.1 31.2 $(\uparrow 1.1)$ 30.3 $(\uparrow 0.2)$ 0.80 0.70 $(\downarrow 0.1)$ 0.50 $(\downarrow 0.3)$
SR52.5 52.7 $(\uparrow 0.2)$ 52.6 $(\uparrow 0.1)$ 43.6 44.0 $(\uparrow 0.4)$ 40.6 $(\downarrow 3.4)$ 4.20 4.30 $(\uparrow 0.1)$ 3.90 $(\uparrow 0.3)$
TruthfulQA MC1 30.4 30.7 ($\uparrow 0.3$) / 34.3 35.1 ($\uparrow 0.8$) / 55.8 56.4 ($\uparrow 0.6$) /
MC2 48.2 48.2 ($=0.0$) / 52.9 53.6 ($\uparrow 0.7$) / 72.1 72.8 ($\uparrow 0.7$) /
@@ 4.2 Bias Evaluation
4.3 Generalisation across Various LLMs
We show the results on TruthfulQA and BBQ of applying ICL, Linear SEA and $\Phi$-SEA on six LLMs with different model families and sizes in Table 3. Specifically, we test them on the LLaMA-2 family [40], Gemma family [39] and Mistral 7B [17], which represent the state-of-the-art open-source LLMs. Linear SEA shows generalisable improvements across all tested LLMs on TruthfulQA. On BBQ, we observe a general trend across different LLMs that linear SEA can marginally improve accuracy, but the other two metrics are mixed. $\Phi$-SEA shows promising performance and can gain increased accuracy, lower unknown-answer rate and stereotypical response rate across all LLMs, except a negligible higher stereotypical response rate for Gemma-it-2B.
@@ 4.5 Post-Editing Performance on Control Tasks
Table 4: Performance of LLaMA-2-Chat-7B and its three SEA edited models for truthfulness and fairness on six control tasks covering multi-task ability, commonsense question answering, and mathematical ability. Details of evaluation are provided in Appendix H.4.
@@
w/ $\Phi$-SEA-Fair 51.93 $\downarrow$ 5.85 14.90 $\downarrow$ 7.92 20.92 $\downarrow$ 1.52 30.05 $\downarrow$ 1.64 45.30 $\downarrow$ 1.15 56.38 $\uparrow$ 5.21
We then leverage six additional control tasks to analyse the editing effects of SEA on other LLM capabilities in Table 4. These tasks include MMLU [14], which serves as the most general benchmark. We then use HellaSwag [19] and Natural Questions [43] to assess the model’s commonsense reasoning and question answering. We use GSM8K [8] and MathQA [1] to verify the model’s ability to solve mathematical tasks. We rely on ToxiGen [13] for assessing the model’s toxicity.
Editing a model’s activations has only a limited impact on other capabilities. We first note that linear editing almost does not hurt the model’s other capabilities (e.g., in commonsense and maths). The non-linear editing causes a small drop in performance in maths tasks, but we observe a more visible decrease in common-sense QA. We attribute this to the limitation mentioned in §2.5, namely that the non-linear projection using feature functions is not theoretically lossless. Qualitative examples in Appendix G also show the high quality and fluency in outputs of SEA-edited models.
4.6 Generalisation of Editing Effects to Similar Tasks
Table 5: Results of all SEA methods editing with BBQ’s demonstrations on all bias categories of CrowS-Pairs. Abbreviations: Aut – Autre, Dis – Disability, Gen – Gender, Nat – Nationality, App – Appearance, Rel – Religion, R/C – Race/Color, SE – Socioeconomic, and SO – Sexual Orientation.
Age Aut Dis Gen Nat App R/C Rel SO SE Avg.
LLaMA-2 75.82 72.73 73.85 61.56 61.11 72.22 53.15 75.68 86.02 71.58 64.16
w/ Linear-SEA 74.73 72.73 72.31 62.19 60.19 70.83 53.35 75.68 86.02 72.11 64.10
w/ $\Phi$-SEA 78.02 72.73 67.69 59.06 52.31 70.83 45.47 67.57 77.42 62.11 57.96
SEA’s editing effect can be generalised to other similar tasks. Finally, we observe that the post-editing improvements can be generalised across new tasks, provided they share some similarities. For example, we observe a strong improvement of the 7B LLaMA-2-Chat model in ToxiGen in Table 4, after projecting the activations towards the less biased directions.
We further assess all SEA methods using BBQ’s demonstrations on the CrowS-Pairs dataset [26], which evaluates a model’s propensity to generate biased outputs (Table 5). We report the percentage of stereotypical sentences rated as more likely than non-stereotypical ones, where a lower percentage indicates less bias. Results demonstrate that both SEA variants effectively reduce model bias across most categories, with $\Phi$-SEA notably decreasing the generation of stereotypical sentences by 7%.
@@ 4.7 Spectral Analysis for the Source of Model Behaviours
We now present an analysis to substantiate why editing the top-$L$ layers proves to be more precise compared to other editing schemes, such as the bottom-$L$ layers. We follow [11, 46, 47] which proposed interpreting the signature as the sum of singular values resulting from an SVD of the cross-covariance for two variables, indicating the degree of correlation between them. We extend this to identify which layers’ activations contain the most information regarding the model’s behaviour. The calculation of the signature value is in Appendix A.
@@ 5 Related Work
Modifying the activations of a trained model, thus altering the model’s behaviour [30, 16, 24, 49, 6, 44] or internal knowledge [9], represents a lightweight method to control the model’s generation. Li et al. [21] probes LLM’s attention heads which are accountable for hallucinations, then edits activations toward truthful directions. Another way is extracting latent vectors directly from the trained model and leveraging these vectors to regulate the model’s inference [41, 37, 32, 49]. Recently, Singh et al. [35] demonstrated the efficacy of fitting an optimal transport from negative to positive activations to facilitate effective non-linear editing. Activation editing finds application in decoding as well, either by contrasting activations from various layers [7] or by using a weaker model to edit activations from a stronger model [22, 45]. Distinct with previous works that use probing [21], contrasting activations [6, 49, 45, 22], or optimal transfer [35], we use the covariance information to find the editing directions for LLM’s activations.